Find or calculate the acceleration of drag car that goes from 0 to 300km/h in 0.56km.
1
Expert's answer
2019-10-24T09:28:55-0400

We have:


"v^2-0^2=2as"

"\left(\frac{300}{3.6}\right)^2-0^2=2(560)a"

"a=6.2\frac{m}{s^2}"